# DAmodel: Hierarchical Bayesian Modelling of DA White Dwarfs for Spectrophotometric Calibration

### v1

@@ 3.5 Priors
Table 1: White dwarf and dust parameters inferred using DAmodel. Parameters for stars with an * were inferred from STIS UV spectra as well as ground based optical spectra. Results shown are the averages and standard deviations from our chains. Note that the achromatic offset $\mu$ cannot be treated exactly as distance modulus, as as the absolute magnitude of each WD is not known without imposing a mass-radius-luminosity relation.
Object $T_{\text{eff}}$ $\log g$ $A_{V}$ $R_{V}$ $\mu$
K mag mag
G191B2B $61941\pm 519$ $7.67\pm 0.03$ $0.0\pm 0$ 0.0 $3.1\pm 0.18$ $52.59\pm 0.01$
GD153 $41437\pm 276$ $7.76\pm 0.03$ $0.02\pm 0.0$ $3.18\pm 0.15$ $53.74\pm 0.01$
GD71 $33719\pm 118$ $7.88\pm 0.01$ $0.02\pm 0.0$ $3.24\pm 0.16$ $53.13\pm 0.01$
WDFS0103-00 $64833\pm 3796$ $7.55\pm 0.06$ $0.13\pm 0.01$ $2.94\pm 0.16$ $60.06\pm 0.05$
WDFS0228-08 $22597\pm 483$ $7.96\pm 0.03$ $0.14\pm 0.02$ $3.11\pm 0.17$ $59.17\pm 0.02$
WDFS0248+33 $33646\pm 483$ $7.29\pm 0.06$ $0.34\pm 0.01$ $3.36\pm 0.08$ $58.32\pm 0.02$
WDFS0727+32 $54754\pm 3101$ $7.79\pm 0.05$ $0.16\pm 0.01$ $2.91\pm 0.13$ $58.77\pm 0.05$
WDFS0815+07 $36439\pm 1623$ $7.22\pm 0.06$ $0.09\pm 0.02$ $3.05\pm 0.18$ $60.1\pm 0.05$
WDFS1024-00 $36012\pm 1350$ $7.79\pm 0.13$ $0.23\pm 0.02$ $3.06\pm 0.15$ $59.09\pm 0.05$
WDFS1110-17 $52672\pm 1921$ $7.8\pm 0.04$ $0.19\pm 0.01$ $3.25\pm 0.11$ $58.56\pm 0.03$
WDFS1111+39 $56759\pm 3481$ $7.84\pm 0.07$ $0.02\pm 0.01$ $3.1\pm 0.18$ $59.39\pm 0.05$
WDFS1206+02 $24243\pm 272$ $7.96\pm 0.02$ $0.08\pm 0.01$ $3.12\pm 0.16$ $58.22\pm 0.01$
WDFS1214+45 $36003\pm 438$ $7.97\pm 0.03$ $0.05\pm 0.01$ $3.11\pm 0.16$ $58.18\pm 0.01$
WDFS1302+10 $46154\pm 887$ $7.87\pm 0.02$ $0.1\pm 0.01$ $2.87\pm 0.11$ $57.71\pm 0.02$
WDFS1314-03 $43685\pm 2696$ $7.79\pm 0.07$ $0.11\pm 0.02$ $3.24\pm 0.16$ $59.69\pm 0.05$
WDFS1514+00 $29853\pm 74$ $7.88\pm 0.01$ $0.18\pm 0.0$ $3.45\pm 0.08$ $55.6\pm 0.01$
WDFS1557+55 $66877\pm 1942$ $7.57\pm 0.05$ $0.05\pm 0.01$ $2.95\pm 0.17$ $58.57\pm 0.02$
WDFS1638+00 $65645\pm 3408$ $7.63\pm 0.13$ $0.23\pm 0.01$ $3.18\pm 0.14$ $59.7\pm 0.04$
WDFS1814+78 $32019\pm 144$ $7.91\pm 0.02$ $0.05\pm 0.0$ $3.06\pm 0.14$ $56.75\pm 0.01$
WDFS2101-05 $29696\pm 358$ $7.86\pm 0.05$ $0.17\pm 0.01$ $3.22\pm 0.13$ $58.54\pm 0.02$
WDFS2329+00 $21729\pm 332$ $7.99\pm 0.03$ $0.19\pm 0.02$ $3.13\pm 0.12$ $57.36\pm 0.02$
WDFS2351+37 $47766\pm 1354$ $7.93\pm 0.06$ $0.37\pm 0.01$ $3.12\pm 0.07$ $58.45\pm 0.03$
WDFS0122-30 $36474\pm 628$ $7.66\pm 0.05$ $0.08\pm 0.01$ $3.08\pm 0.16$ $58.86\pm 0.02$
WDFS0238-36 $24094\pm 225$ $7.88\pm 0.04$ $0.21\pm 0.01$ $3.13\pm 0.1$ $57.48\pm 0.01$
WDFS0458-56 $31128\pm 218$ $7.79\pm 0.04$ $0.04\pm 0.01$ $3.1\pm 0.17$ $57.9\pm 0.01$
WDFS0541-19 $21144\pm 195$ $7.88\pm 0.03$ $0.1\pm 0.01$ $3.25\pm 0.16$ $57.54\pm 0.01$
WDFS0639-57 $54352\pm 1811$ $7.71\pm 0.07$ $0.19\pm 0.01$ $3.61\pm 0.12$ $58.92\pm 0.03$
WDFS0956-38 $19791\pm 79$ $7.87\pm 0.01$ $0.12\pm 0.01$ $3.38\pm 0.14$ $56.97\pm 0.01$
WDFS1055-36 $30682\pm 199$ $7.93\pm 0.03$ $0.15\pm 0.01$ $3.31\pm 0.11$ $58.0\pm 0.01$
WDFS1434-28 $20790\pm 128$ $7.87\pm 0.03$ $0.22\pm 0.01$ $3.33\pm 0.11$ $57.07\pm 0.01$
WDFS1837-70 $19975\pm 91$ $7.89\pm 0.02$ $0.14\pm 0.01$ $3.02\pm 0.12$ $56.88\pm 0.01$
WDFS1930-52 $38498\pm 632$ $7.79\pm 0.04$ $0.17\pm 0.01$ $3.33\pm 0.11$ $57.86\pm 0.02$
WDFS1206-27 $35884\pm 361$ $7.8\pm 0.04$ $0.15\pm 0.01$ $3.35\pm 0.11$ $56.77\pm 0.01$
WDFS1535-77 $62316\pm 1428$ $8.97\pm 0.03$ $0.07\pm 0.0$ $2.85\pm 0.15$ $57.54\pm 0.02$
WDFS2317-29 $25063\pm 224$ $7.82\pm 0.03$ $0.08\pm 0.01$ $2.83\pm 0.16$ $57.97\pm 0.01$
@@ 4 Results and Discussion
In this section we present our calibration results using our hierarchical Bayesian model. The main results presented are those that use all the optical ground-based spectra, the STIS UV space-based spectra for the 19 WDs that have coverage, and the six HST/WFC3 bands ranging the UV to the NIR. The published results are also those that infer $R^{s}_{V}$ for the WDs, rather than keeping it constant as in previous analyses. The prior used in the $R^{s}_{V}$ inference is a truncated normal prior $R^{s}_{V}\sim TN(3.1,0.18,1.2,\infty)$ motivated by Schlafly et al. (2016), where $\sim$ denotes “is sampled from”. In these main results we do not infer dust population parameters to avoid the effects of shrinkage, where inferences of object-level parameters are skewed towards the inferred population average. Our results are on the CALSPEC Bohlin et al. (2020) magnitude system, tied at the F475W band, but may easily be converted to the Bohlin (2014) system using the zeropoints presented in Table 4. The SEDs for all 35 WDs in the network are defined across the entire range of the Gordon et al. (2023) dust relation, spanning from 912 Å to 32 $\mu$m, and are publicly accessible at: https://zenodo.org/records/14339961
@@ 4.2 Spectroscopic Spline Residuals
Table 2: Inferred HST/WFC3 photometric systematics obtained in our final results.
@@
Table 3: HST/WFC3 synthetic photometry and residuals from our final results. Units are in AB magnitudes and photometry for stars with an * was inferred using STIS UV spectra as well as ground-based optical spectra.
@@ 4.3 HST/WFC3 Photometric Residuals and Systematics
In our work we achieve the lowest residuals yet when using our modelled SEDs to predict the observed data, illustrated in Figure 7. We use the 2000 samples of WD and dust parameters to forward model 2000 randomly sampled SEDs per object. To obtain the synthetic photometry we integrate the SED samples using Equation (3) over the six HST/WFC3 bandpasses described in Calamida et al. (2019, 2022b). We then define our synthetic photometry and uncertainty as the mean and standard deviation of the 2000 sampled magnitudes in each band for each object. We use the average of each inferred systematic (presented in Table 2) to correct for instrumental effects to derive the calibrated observed photometry. A detailed breakdown of the residual and synthetic photometry of each object is shown in Table 3.
@@
Figure 9: Photometric residuals when comparing our synthetic DES photometry with real DES observations (Abbott et al., 2021) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes.
@@
An important systematic to consider is the count-rate nonlinearity in the F160W band. In our analysis, we infer the CRNL to be $\alpha_{\text{F160W}}$ = -3.19 $\pm$ 0.31 mmag/mag. This constraint is within $1\sigma$ of both the published CRNL value of -3.12 $\pm$ 0.32 mmag/mag from Bohlin & Deustua (2019) and the combined result of -3.0 $\pm$ 0.24 mmag/mag from Riess et al. (2019). The result is also close to the post hoc result of -2.6$\pm$0.52 mmag/mag from Narayan et al. (2019). The constraint, however, is in disagreement with the result of -1.74 $\pm$ 0.32 mmag/mag from Axelrod et al. (2023), that used the same WD sample as this work. Interestingly, if we change our modelling decisions to match those of Axelrod et al. (2023), by using Fitzpatrick (1999), a constant $R_{V}=3.1$ and ignoring the STIS UV data, we obtain $\alpha_{\text{F160W}}$ = -2.12 $\pm$ 0.30 mmag/mag. This result is a strong suggestion that the inferred value of the NIR CRNL is sensitive to dust modelling assumptions. In inference, our model is determining which effects are from natural sources such as dust and which are from instrumental systematics. If we change the definition of what is physically possible, such as fixing or changing dust relations, it will then impact the extent to which effects are inferred to be coming from systematics. The inferred CRNL’s dependency on dust is also supported by correlations shown in Figure 17.
@@ 4.4 Photometric Residuals with other surveys
Figure 9 shows the residual scatter for seven of our DA WDs in DES DR2 that were observed with DECam (Abbott et al., 2021). We see residual bias in $g$, $r$ and $i$ bands between 12 and 13 mmag, which then reduces to 4 mmag in the $z$ band. We show corresponding results in Figure 21 using SDSS (Holberg & Bergeron, 2006), Figure 22 using PS1 DR2 (Flewelling et al., 2020), Figure 23 using Gaia DR3 (Gaia Collaboration et al., 2023) and Figure 24 using DECaLS (Schlegel et al., 2021) data. Residual bias with other surveys is highly dependent on the magnitude systems used in calibration of both our work and the surveys themselves. In our method we set the primary standard magnitudes as constants in one band according to Bohlin et al. (2020) CALSPEC. We find that when we set the magnitudes to constant in all six bands according to Bohlin (2014) CALSPEC, residual scatter with other surveys decreases in the optical but increases towards the NIR.

### v2

@@ 3.5 Priors
Table 1: White dwarf and dust parameters inferred using DAmodel. Parameters for stars with an * were inferred from STIS UV spectra as well as ground based optical spectra. Results shown are the averages and standard deviations from our chains. Note that the achromatic offset $\mu$ cannot be treated exactly as distance modulus, as as the absolute magnitude of each WD is not known without imposing a mass-radius-luminosity relation. This table is available in the online supplementary material.
Object R.A. Dec. $T_{\text{eff}}$ $\log g$ $A_{V}$ $R_{V}$ $\mu$
hh:mm:ss.s dd:mm:ss.s K mag mag
G191B2B 05:05:30.61 +52:49:51.96 $61941\pm 519$ $7.67\pm 0.03$ $0.0\pm 0$ 0.0 $3.1\pm 0.18$ $52.59\pm 0.01$
GD153 12:57:02.34 +22:01:52.68 $41437\pm 276$ $7.76\pm 0.03$ $0.02\pm 0.0$ $3.18\pm 0.15$ $53.74\pm 0.01$
GD71 05:52:27.61 +15:53:13.75 $33719\pm 118$ $7.88\pm 0.01$ $0.02\pm 0.0$ $3.24\pm 0.16$ $53.13\pm 0.01$
WDFS0103-00 01:03:22.201 -00:20:47.800 $64833\pm 3796$ $7.55\pm 0.06$ $0.13\pm 0.01$ $2.94\pm 0.16$ $60.06\pm 0.05$
WDFS0122-30 01:22:00.725 -30:52:03.95 $36474\pm 628$ $7.66\pm 0.05$ $0.08\pm 0.01$ $3.08\pm 0.16$ $58.86\pm 0.02$
WDFS0228-08 02:28:17.183 -08:27:16.301 $22597\pm 483$ $7.96\pm 0.03$ $0.14\pm 0.02$ $3.11\pm 0.17$ $59.17\pm 0.02$
WDFS0238-36 02:38:24.969 -36:02:23.222 $24094\pm 225$ $7.88\pm 0.04$ $0.21\pm 0.01$ $3.13\pm 0.1$ $57.48\pm 0.01$
WDFS0248+33 02:48:54.965 33:45:48.244 $33646\pm 483$ $7.29\pm 0.06$ $0.34\pm 0.01$ $3.36\pm 0.08$ $58.32\pm 0.02$
WDFS0458-56 04:58:23.133 -56:37:33.434 $31128\pm 218$ $7.79\pm 0.04$ $0.04\pm 0.01$ $3.1\pm 0.17$ $57.9\pm 0.01$
WDFS0541-19 05:41:14.759 -19:30:38.896 $21144\pm 195$ $7.88\pm 0.03$ $0.1\pm 0.01$ $3.25\pm 0.16$ $57.54\pm 0.01$
WDFS0639-57 06:39:41.468 -57:12:31.164 $54352\pm 1811$ $7.71\pm 0.07$ $0.19\pm 0.01$ $3.61\pm 0.12$ $58.92\pm 0.03$
WDFS0727+32 07:27:52.752 32:14:16.046 $54754\pm 3101$ $7.79\pm 0.05$ $0.16\pm 0.01$ $2.91\pm 0.13$ $58.77\pm 0.05$
WDFS0815+07 08:15:08.782 07:31:45.775 $36439\pm 1623$ $7.22\pm 0.06$ $0.09\pm 0.02$ $3.05\pm 0.18$ $60.1\pm 0.05$
WDFS0956-38 09:56:57.009 -38:41:30.269 $19791\pm 79$ $7.87\pm 0.01$ $0.12\pm 0.01$ $3.38\pm 0.14$ $56.97\pm 0.01$
WDFS1024-00 10:24:30.912 -00:32:07.16 $36012\pm 1350$ $7.79\pm 0.13$ $0.23\pm 0.02$ $3.06\pm 0.15$ $59.09\pm 0.05$
WDFS1055-36 10:55:25.356 -36:12:14.731 $30682\pm 199$ $7.93\pm 0.03$ $0.15\pm 0.01$ $3.31\pm 0.11$ $58.0\pm 0.01$
WDFS1110-17 11:10:59.436 -17:09:54.308 $52672\pm 1921$ $7.8\pm 0.04$ $0.19\pm 0.01$ $3.25\pm 0.11$ $58.56\pm 0.03$
WDFS1111+39 11:11:27.313 39:56:28.105 $56759\pm 3481$ $7.84\pm 0.07$ $0.02\pm 0.01$ $3.1\pm 0.18$ $59.39\pm 0.05$
WDFS1206+02 12:06:50.41 02:01:42.138 $24243\pm 272$ $7.96\pm 0.02$ $0.08\pm 0.01$ $3.12\pm 0.16$ $58.22\pm 0.01$
WDFS1206-27 12:06:20.354 -27:29:40.639 $35884\pm 361$ $7.8\pm 0.04$ $0.15\pm 0.01$ $3.35\pm 0.11$ $56.77\pm 0.01$
WDFS1214+45 12:14:05.111 45:38:18.626 $36003\pm 438$ $7.97\pm 0.03$ $0.05\pm 0.01$ $3.11\pm 0.16$ $58.18\pm 0.01$
WDFS1302+10 13:02:34.422 10:12:38.717 $46154\pm 887$ $7.87\pm 0.02$ $0.1\pm 0.01$ $2.87\pm 0.11$ $57.71\pm 0.02$
WDFS1314-03 13:14:45.046 -03:14:15.685 $43685\pm 2696$ $7.79\pm 0.07$ $0.11\pm 0.02$ $3.24\pm 0.16$ $59.69\pm 0.05$
WDFS1434-28 14:34:59.528 -28:19:03.295 $20790\pm 128$ $7.87\pm 0.03$ $0.22\pm 0.01$ $3.33\pm 0.11$ $57.07\pm 0.01$
WDFS1514+00 15:14:21.277 00:47:52.380 $29853\pm 74$ $7.88\pm 0.01$ $0.18\pm 0.0$ $3.45\pm 0.08$ $55.6\pm 0.01$
WDFS1535-77 15:35:45.179 -77:24:44.832 $62316\pm 1428$ $8.97\pm 0.03$ $0.07\pm 0.0$ $2.85\pm 0.15$ $57.54\pm 0.02$
WDFS1557+55 15:57:45.38 55:46:09.361 $66877\pm 1942$ $7.57\pm 0.05$ $0.05\pm 0.01$ $2.95\pm 0.17$ $58.57\pm 0.02$
WDFS1638+00 16:38:00.352 00:47:17.739 $65645\pm 3408$ $7.63\pm 0.13$ $0.23\pm 0.01$ $3.18\pm 0.14$ $59.7\pm 0.04$
WDFS1814+78 18:14:24.078 78:54:03.084 $32019\pm 144$ $7.91\pm 0.02$ $0.05\pm 0.0$ $3.06\pm 0.14$ $56.75\pm 0.01$
WDFS1837-70 18:37:17.906 -70:02:52.513 $19975\pm 91$ $7.89\pm 0.02$ $0.14\pm 0.01$ $3.02\pm 0.12$ $56.88\pm 0.01$
WDFS1930-52 19:30:18.995 -52:03:46.55 $38498\pm 632$ $7.79\pm 0.04$ $0.17\pm 0.01$ $3.33\pm 0.11$ $57.86\pm 0.02$
WDFS2101-05 21:01:50.667 -05:45:51.159 $29696\pm 358$ $7.86\pm 0.05$ $0.17\pm 0.01$ $3.22\pm 0.13$ $58.54\pm 0.02$
WDFS2317-29 23:17:20.294 -29:03:21.647 $25063\pm 224$ $7.82\pm 0.03$ $0.08\pm 0.01$ $2.83\pm 0.16$ $57.97\pm 0.01$
WDFS2329+00 23:29:41.321 00:11:07.565 $21729\pm 332$ $7.99\pm 0.03$ $0.19\pm 0.02$ $3.13\pm 0.12$ $57.36\pm 0.02$
WDFS2351+37 23:51:44.274 37:55:42.569 $47766\pm 1354$ $7.93\pm 0.06$ $0.37\pm 0.01$ $3.12\pm 0.07$ $58.45\pm 0.03$
Notes: Coordinates are from Gaia DR3 at epoch J2016.0 (Gaia Collaboration et al., 2023). Corresponding proper motions for each faint standard in our network can be found in Table 2 of Axelrod et al. (2023).
@@ 4 Results and Discussion
In this section we present our calibration results using our hierarchical Bayesian model. The main results presented are those that use all the optical ground-based spectra, the STIS UV space-based spectra for the 19 WDs that have coverage, and the six HST/WFC3 bands ranging the UV to the NIR. The published results are also those that infer $R^{s}_{V}$ for the WDs, rather than keeping it constant as in previous analyses. The prior used in the $R^{s}_{V}$ inference is a truncated normal prior $R^{s}_{V}\sim\mathcal{TN}(3.1,0.18,1.2,\infty)$ motivated by Schlafly et al. (2016), where $\sim$ denotes “is sampled from”. In these main results we do not infer dust population parameters to avoid the effects of shrinkage, where inferences of object-level parameters are skewed towards the inferred population average. Our results are on the CALSPEC Bohlin et al. (2020) magnitude system, tied at the F475W band, but may easily be converted to the Bohlin (2014) system using the zeropoints presented in Table 5. The SEDs for all 35 WDs in the network are defined across the entire range of the Gordon et al. (2023) dust relation, spanning from 912 Å to 32 $\mu$m, and are publicly accessible at: https://zenodo.org/records/14339960
@@ 4.2 Spectroscopic Spline Residuals
Table 2: Inferred HST/WFC3 photometric systematics obtained in our final results. This table is available in the online supplementary material.
@@
Table 3: HST/WFC3 synthetic photometry and residuals (labelled Synth. and Resid. respectively) from our final results. Units are in AB magnitudes and photometry for stars with an * was inferred using STIS UV spectra as well as ground-based optical spectra. This table is available in the online supplementary material.
@@ 4.3 HST/WFC3 Photometric Residuals and Systematics
In our work we achieve the lowest residuals yet when using our modelled SEDs to predict the observed data, illustrated in Figure 7. We use the 2000 samples (500 posterior samples from each of the four chains) of WD and dust parameters to forward model 2000 randomly sampled SEDs per object. To obtain the synthetic photometry we integrate the SED samples using Equation (3) over the six HST/WFC3 bandpasses described in Calamida et al. (2019, 2022b). We then define our synthetic photometry and uncertainty as the mean and standard deviation of the 2000 sampled magnitudes in each band for each object. We use the average of each inferred systematic (presented in Table 2) to correct for instrumental effects to derive the calibrated observed photometry. A detailed breakdown of the residual and synthetic photometry of each object is shown in Table 3.
@@
Figure 9: Photometric residuals when comparing our synthetic DES photometry with real DES observations (Abbott et al., 2021) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes. Plotted in white are the residuals when using the Bohlin (2014) CALSPEC system rather than the Bohlin et al. (2020) system.
@@
An important systematic to consider is the CRNL in the F160W band. In our analysis, we infer the CRNL to be $\alpha_{\text{F160W}}$ = -3.19 $\pm$ 0.31 mmag/mag. This constraint is within $1\sigma$ of both the published CRNL value of -3.12 $\pm$ 0.32 mmag/mag from Bohlin & Deustua (2019) and the combined result of -3.0 $\pm$ 0.24 mmag/mag from Riess et al. (2019). The result is also close to the post hoc result of -2.6$\pm$0.52 mmag/mag from Narayan et al. (2019). The constraint, however, is in disagreement with the result of -1.74 $\pm$ 0.32 mmag/mag from Axelrod et al. (2023), that used the same WD sample as this work. Interestingly, if we change our modelling decisions to match those of Axelrod et al. (2023), by using Fitzpatrick (1999), a constant $R_{V}=3.1$ and ignoring the STIS UV data, we obtain $\alpha_{\text{F160W}}$ = -2.12 $\pm$ 0.30 mmag/mag. This result is a strong suggestion that the inferred value of the NIR CRNL is sensitive to dust modelling assumptions. In inference, our model is determining which effects are from natural sources such as dust and which are from instrumental systematics. If we change the definition of what is physically possible, such as fixing or changing dust relations, it will then impact the extent to which effects are inferred to be coming from systematics. The inferred CRNL’s dependency on dust is also supported by correlations shown in Figure 19.
@@ 4.4 Photometric Residuals with other surveys
Figure 9 shows the residual scatter for seven of our DA WDs in DES DR2 that were observed with DECam (Abbott et al., 2021). We see residual bias in $g$, $r$ and $i$ bands between 12 and 13 mmag, which then reduces to 4 mmag in the $z$ band. We show corresponding results in Figure 23 using SDSS (Holberg & Bergeron, 2006), Figure 24 using PS1 DR2 (Flewelling et al., 2020), Figure 25 using Gaia DR3 (Gaia Collaboration et al., 2023) and Figure 26 using DECaLS (Schlegel et al., 2021) data. Residual bias with other surveys is highly dependent on the magnitude systems used in calibration of both our work and the surveys themselves. In our method we set the primary standard magnitudes as constants in one band according to Bohlin et al. (2020) CALSPEC. The other surveys that we calculate residuals with have been calibrated with respect to the CALSPEC Bohlin (2014) system or older. In Figure 9 and the Figures in Appendix 4.4 we also show the residuals if we were to use the CALSPEC Bohlin (2014) system. These results are comparable to those in Axelrod et al. (2023). This change to the older system dramatically reduces the size residuals, particularly with Gaia in Figure 25, where we have an average residual of -1 mmag in the $G$ band when using the Bohlin (2014) system and the same scaled definition of Vega (Riello et al., 2021). Despite getting better agreement with other surveys when using the older CALSPEC system, we maintain our decision to publish the network with respect to the more accurate Bohlin et al. (2020) CALSPEC system and encourage other surveys to adopt the same for their absolute calibration.